40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 396     Entry time: Sun Mar 23 00:56:42 2008
Author: John 
Type: Update 
Category: LSC 
Subject: More on 3f  
We ended our last attempt at 3f locking concerned about the beam size on PD6. I investigated tonight. The beam was not obviously overfilling the diode and a quick tweak of the steering mirror revealed a decent plateaux. Nevertheless we decided to try a different approach to see if we found the same problems as before on a different diode.

This time our 3f diode was Refl 33. I put a splitter on the output of the diode at the LSC rack sending one half into the usual refl 33 board, the other into refl DD 199 (which is demodulating at 99Mhz).

I got as far as handing off PRC to the 3f signal in lock. More work needed.
ELOG V3.1.3-